OpenStudy (anonymous):

Solve using the guess-and-test strategy. Michael and Susan have a combined age of 17 years. Michael is 5 years older than Susan. How old is Michael? A. 5 B.6 C.11 D.22 Medal and fan if you help me What I got was is Michal is 13 years old please correct me if I am wrong MEDAL AND FAN

OpenStudy (confluxepic):

Does 13+8=17. I don't think so.

OpenStudy (confluxepic):

Michael is 11. Susan is 6.

OpenStudy (confluxepic):

11+6=17 11-6=5

OpenStudy (anonymous):

thx \

OpenStudy (confluxepic):

You're welcome.

OpenStudy (confluxepic):

\[\huge \color{blue}{:)}\]